OpenStudy (zuzu11):

How many solutions does the equation have? 4(z−5)+2=4z−18

OpenStudy (studygurl14):

First simplify 4(z - 5) + 2 = 4z - 18 4z - 20 + 2 = 4z - 18 The 4z s on both sides cancel out -20 + 2 = -18 -18 = -18 Since this works out and there are no more variables, there is an infinite number of solutions

OpenStudy (zuzu11):

so the answer is -18

OpenStudy (zuzu11):

oh wait nvm the answer is many solutions, thanks!! :)

OpenStudy (precal):

both sides are the same, therefore many solutions 4z-18=4z-18